Prove the following:


i. [AB,C] = [Â,C]B + Â[B,C]


ii. ihq = qH - Hq given that H = mq


1
Expert's answer
2021-08-02T08:40:08-0400

(i)

"[AB,C]=ABC-CAB"

"=ABC-ACB+ACB-CAB"

"=A(BC-CB)+(AC-CA)B"

"=A[B,C]+[A,C]B"

(ii)

"i\hbar \dot q=[q,H]=qH-Hq=q(mq)-(mq)q=0"